L2kool:

14n+6p-8n= use 18p for n

Astrid1:

Alright so here's our equation provided: 14n+6p-8n= use 18p for n Group the like terms. 6p+14n-8n= 18p 6p+6n=18p Now you'll want to subtract 6p from both sides. 6p+6n-6p=18p-6p 6n=12p Now you'll want to divide. We'll take 6n and divide it by 6. WHich obviously =0 And so we're going to divide 12p by 6. Whats 6 divided by 12?

Astrid1:

Make sure to add the variable when you get your answer. It should be 2p.

Astrid1:

If I get this wrong, I apologize, it has been awhile since I've done equations like this.